OpenStudy (anonymous):

Let f(x)= sqrt of x and g(x)=cubed root of 1-x. Find a formula for g(f(x) and find it's domain

OpenStudy (danjs):

\[g(f(x))=\sqrt[3]{1-f(x)} = \sqrt[.3]{1-\sqrt{x}}\]

OpenStudy (dinamix):

\[\sqrt[3]{1-\sqrt{x}}\]@Blairortiz

OpenStudy (dinamix):

lol @DanJS

OpenStudy (danjs):

hit post at the same time

OpenStudy (dinamix):

yeah

OpenStudy (danjs):

@Blairortiz Do you know the domain of the square root and the cubed root functions?

OpenStudy (anonymous):

No

OpenStudy (dinamix):

\[x ^{1/2} =\sqrt{x}\]

OpenStudy (dinamix):

\[\sqrt[3]{x}=x ^{3/2}\]

OpenStudy (anonymous):

So the domain is x^3/2 and x^1/2?

OpenStudy (danjs):

\[\sqrt x\] For this x has to be greater than or equal to zero, so the domain of f(x) is all x greater than or equal to zero...

OpenStudy (danjs):

The cubed root function can have any real number as a value of x...

OpenStudy (danjs):

So for the g(f(x)), the domain is all x greater than or equal to zero, the red function in the graph